A greenhouse gas regulation will cost $15 billion in year 0. The benefits generated from the regulation will be spread over five years with annual benefits indicated in the table below.

a. Should this regulation be adopted if the annual an interest rate is 5% per year.

b. a. What is the internal rate of return for this regulation?

Year: 1, 2,  3, 4, 5

Benefit (Billions):$1, $3, $3, $5, $8

Hint: The discount factor in year t is given as [1/(1+r)]^t where r is the annual rate of interest expressed as a fraction. I recommend using a computer to solve part b.
Answer #1

(a)

NPV ($ Billion) = - 15 + 1 x P/F(5%, 1) + 3 x P/F(5%, 2) + 3 x P/F(5%, 3) + 5 x P/F(5%, 4) + 8 x P/F(5%, 5)

= - 15 + 1 x 0.95 + 3 x 0.91 + 3 x 0.86 + 5 x 0.84 + 8 x 0.80

= - 15 + 0.95 + 2.73 + 2.58 + 4.2 + 6.4

= 1.86

Since NPV > 0, the regulation should be adopted.

(b)

IRR is computed using excel IRR function, as follows.

Year Cash Flow ($ Billion)
0 -15
1 1
2 3
3 3
4 5
5 8
IRR = 8.00%